# MMSR: Symbolic Regression is a Multimodal Task

### 10 Mar 2024 v2

@@ Abstract
Mathematical formulas are the crystallization of human wisdom in exploring the laws of nature for thousands of years. Describing the complex laws of nature with a concise mathematical formula is a constant pursuit of scientists and a great challenge for artificial intelligence. This field is called symbolic regression. Symbolic regression was originally formulated as a combinatorial optimization problem, and GP and reinforcement learning algorithms were used to solve it. However, GP is sensitive to hyperparameters, and these two types of algorithms are inefficient. To solve this problem, researchers treat the mapping from data to expressions as a translation problem. And the corresponding large-scale pre-trained model is introduced. However, the data and expression skeletons do not have very clear word correspondences as the two languages do. Instead, they are more like two modalities (e.g., image and text). Therefore, in this paper, we proposed MMSR. The SR problem is solved as a pure multimodal problem, and contrastive learning is also introduced in the training process for modal alignment to facilitate later modal feature fusion. It is worth noting that in order to better promote the modal feature fusion, we adopt the strategy of training contrastive learning loss and other losses at the same time, which only needs one-step training, instead of training contrastive learning loss first and then training other losses. Because our experiments prove training together can make the feature extraction module and feature fusion module running-in better. Experimental results show that compared with multiple large-scale pre-training baselines, MMSR achieves the most advanced results on multiple mainstream datasets including SRBench.
@@ 1 Introduction
The large-scale pre-trained model gained fast search speed because once the model finished training, we only needed a single forward pass to get the expressions. However, they ignore the different data formats of the input and output of SR. That is, previous methods treat the SR as a machine translation task: translate the data points into corresponding expressions. However, it is clear that the input and output of SR are very different from the two languages in translation, and they do not have an obvious word correspondence like languages. They are more like two different modalities (like image and text). Therefore, it is perhaps more reasonable to think of SR as a multimodal task.
To tackle the problems of two different modalities, we use a modal alignment module in our Transformer to make the encoded data feature close to the skeleton feature. We name our model as MMSR. Specifically, we use the SetTransformer (Lee et al., 2019) as a data encoder for it presents permutation invariance. And, decodes the skeleton of expression with a causal masking transformer decoder. Unlike standard decoder transformers, MMSR omits cross-attention in the first half of the decoder layers to encode unimodal skeleton representations, and cascades the rest of the decoder layers, cross-attending to the data encoder for multimodal data-skeleton representations. We use contrastive learning (Chuang et al., 2020) to align the modal features, so that the matching data features and skeleton features are as close as possible in the feature space, and the non-matching features are as far away from each other as possible. Additionally, we train the model with two extra losses: the cross-entropy loss used for predicting the token of expression skeletons. Second, one of the limitations of pre-train models is that they focus on the supervised pretraining goals borrowed from symbol generation, i.e., they are trained solely with the token-level cross-entropy (CE) loss, which can result in equations that may exhibit high token-level similarities but are suboptimal with respect to equation-specific objectives such as fitting accuracy(Shojaee et al., 2024). To solve this problem, we directly predict the numerical value of the constants and use MSE to measure the loss between the predicted constants and the true constants in the skeletons. We summarize our contributions as follows:
@@ 3.1 Architecture
As depicted in Figure 2, the MMSR consists of two transformer-based encoders, each tailored for learning the symbolic or data representations of mathematical functions. Except that, a decoder receives the encoded numeric representations to decode the features to mathematical expressions. Additionally, an alignment module is designed to match the feature representations of numeric inputs and symbolic inputs. The contrastive learning (Chuang et al., 2020) is used to make the distance between the data and expression skeleton features that are derived from the same expression closer. During training, MMSR receives synthetically created symbolic equations and their associated numeric data as inputs to the symbolic encoder and numeric encoder, respectively. In total, MMSR is trained on approximately 10 million synthetic paired (numeric, symbolic) examples.
@@ 3.1.2 Expression Skeleton Encoder
Symbolic expression by and other basic operators, where [C-5,…,C5] represents a constant placeholder. Each expression can be represented by an expression binary tree, and we can get a symbol sequence by expanding the expression binary tree in preorder. Furthermore, we treat each symbol as a and then embed it.
In particular, to solve the problem that expressions may exhibit high token-level similarities but are suboptimal for equation-specific objectives such as fitting accuracy, which is caused by only using the cross-entropy loss between the generated sequence and the real sequence, we encode the expression symbol sequence and the constant sequence separately. For example, for the expression , we would have the symbolic sequence . And constant sequence . It is worth noting that we replace constants with special symbols. The constants are encoded using a scientific-like notation where a constant C is represented as a tuple of the exponent and the mantissa , . Where belongs to integers between [-5,5] and belongs to [-1,1]. For example, for the constant 0.021 above, we use the placeholder C-1 in the skeleton, where is equal to 0.21 and is equal to -1, so
@@ 3.1.3 Expression Decoder
The decoder autoregressively generates the symbols sequence and the constants sequences given the encoder’s feature. We can think of each token as a token in a text generation task. Then, n candidate expressions are obtained by sampling sequentially using beam search according to the predicted probability distribution of the next symbol. And we’ll get n constants sequences corresponding to n skeletons of expressions. Next, the constant placeholder ‘[C-5,…, C5]’ in the skeleton is replaced by the corresponding constant to obtain the expression with the rough constant. Finally, we fine-tune the rough constant with BFGS to obtain the optimal expression.
@@ 4.3 Ablation Study
To verify whether the introduction of contrastive learning plays a positive role in the algorithm performance, we conduct the ablation experiment as follows. On the premise of ensuring that the training data and other factors are exactly the same, According to Equation 2, we did three sets of experiments, in which and were always equal to 1, and was set to 0.0, 0.1, and 1.0 in turn. The specific experimental results are shown in Table 1.
This can be clearly seen from the table. Other conditions being equal, as the proportion of contrastive learning loss () decreases gradually, the performance of the algorithm also decreases gradually. In addition, the overall performance with the contrastive loss (lambda2 = 1.0 and 0.1) is better than without the contrastive loss (lambda = 0.0).
@@ 4.4 Effects of modality alignment
In order to verify whether the extracted features before and after the introduction of contrastive learning are as expected: Data and skeleton features that match each other are close to each other in the feature space, whereas those that do not match are far away from each other. We first randomly generate 50 expression samples, feed these samples into the model without contrastive learning, and then obtain 50 pairs of features for the data and skeleton. We calculate the cosine distance between the extracted 50 data features and 50 skeleton features and then obtain the 50*50 cosine distance matrix A. Finally, we normalize the matrix A and visualize it to get fig 3(a). Similarly, we feed these samples into a model that uses contrastive learning, resulting in heatmap 3(b).
@@ 4.5 Impact of the size of the training data
To test the impact of pre-training data on model test performance, we train our MMSR model on increasingly larger datasets. More specifically, we use datasets consisting of 10K, 100K, 1M, 5M, and 10M samples to train a model respectively. Except for the amount of data, every other aspect of training is the same as described above, and all are trained for the same epochs.
@@ 4.6 Effect of the introduction method of contrastive learning
Figure.3 shows the comparison plot between the One step strategy and the Tow step strategy. To prove the superiority of our adopted strategy, we conduct comparative experiments on five datasets. From figure 6, we can find that in terms of contrastive learning introduction method, the One-step strategy we adopted is slightly better than the Tow-step strategy in terms of effect. As the size of the dataset increases, this advantage becomes more and more obvious.
We believe that the specific reason for this phenomenon is that in the One-step strategy, contrastive loss and other losses are trained together. In the process of training, the feature extraction module and the feature fusion module are interrelated. This requires the feature extraction module not only to make the extracted data and skeleton features close in the feature space but also to make the extracted features better adapt to the later feature fusion module. So that the overall loss is smaller. In short, the one-step strategy can make each module better fit each other in the process of training together.

### 14 Mar 2024 v4

@@ Abstract
Mathematical formulas are the crystallization of human wisdom in exploring the laws of nature for thousands of years. Describing the complex laws of nature with a concise mathematical formula is a constant pursuit of scientists and a great challenge for artificial intelligence. This field is called symbolic regression. Symbolic regression was originally formulated as a combinatorial optimization problem, and GP and reinforcement learning algorithms were used to solve it. However, GP is sensitive to hyperparameters, and these two types of algorithms are inefficient. To solve this problem, researchers treat the mapping from data to expressions as a translation problem. And the corresponding large-scale pre-trained model is introduced. However, the data and expression skeletons do not have very clear word correspondences as the two languages do. Instead, they are more like two modalities (e.g., image and text). Therefore, in this paper, we proposed MMSR. The SR problem is solved as a pure multimodal problem, and contrastive learning is also introduced in the training process for modal alignment to facilitate later modal feature fusion. It is worth noting that to better promote the modal feature fusion, we adopt the strategy of training contrastive learning loss and other losses at the same time, which only needs one-step training, instead of training contrastive learning loss first and then training other losses. Because our experiments prove training together can make the feature extraction module and feature fusion module wearing-in better. Experimental results show that compared with multiple large-scale pre-training baselines, MMSR achieves the most advanced results on multiple mainstream datasets including SRBench.
@@ 1 Introduction
The large-scale pre-trained model gained fast search speed because once the model finished training, we only needed a single forward pass to get the expressions. However, they ignore the different data formats of the input and output of SR. That is, previous methods treat the SR as a machine translation task: translate the data points into corresponding expressions. However, it is clear that the input and output of SR are very different from the two languages in translation, and they do not have obvious word correspondence like languages. They are more like two different modalities (like image and text). Therefore, it is perhaps more reasonable to think of SR as a multimodal task.
To tackle the problems of two different modalities, we use a modal alignment module in our model to make the encoded data feature close to the skeleton feature. We name our model as MMSR. Specifically, we use the SetTransformer (Lee et al., 2019) as a data encoder for it presents permutation invariance. And, decodes the skeleton of expression with a causal masking transformer decoder. Unlike standard decoder transformers, MMSR omits cross-attention in the first half of the decoder layers to encode unimodal skeleton representations, and cascades the rest of the decoder layers, cross-attending to the data encoder for multimodal data-skeleton representations. We use contrastive learning (Chuang et al., 2020) to align the modal features, so that the matching data features and skeleton features are as close as possible in the feature space, and the non-matching features are as far away from each other as possible. Additionally, we train the model with two extra losses. First, the cross-entropy loss is used for predicting the token of expression skeletons. Second, one of the limitations of pre-train models is that they focus on the supervised pretraining goals borrowed from symbol generation, i.e., they are trained solely with the token-level cross-entropy (CE) loss, which can result in equations that may exhibit high token-level similarities but are suboptimal with respect to equation-specific objectives such as fitting accuracy(Shojaee et al., 2024). To solve this problem, we directly predict the numerical value of the constants and use MSE to measure the loss between the predicted constants and the true constants in the skeletons. We summarize our contributions as follows:
@@ 3.1 Architecture
As depicted in Figure 2, and Algorithm 1 the MMSR consists of two transformer-based encoders, each tailored for learning the symbolic or data representations of mathematical functions. Except that, a decoder receives the encoded numeric representations to decode the features to mathematical expressions. Additionally, an alignment module is designed to match the feature representations of numeric inputs and symbolic inputs. The contrastive learning (Chuang et al., 2020) is used to make the distance between the data and expression skeleton features that are derived from the same expression closer. During training, MMSR receives synthetically created symbolic equations and their associated numeric data as inputs to the symbolic encoder and numeric encoder, respectively. In total, MMSR is trained on approximately 10 million synthetic paired (numeric, symbolic) examples.
@@ 3.1.2 Expression Skeleton Encoder
Symbolic expression by and other basic operators, where [C-5,…,C5] represents a constant placeholder. Each expression can be represented by an expression binary tree, and we can get a symbol sequence by expanding the expression binary tree in preorder, We call this the expression skeleton. Furthermore, we treat each symbol as a and then embed it.
In particular, to solve the problem that expressions may exhibit high token-level similarities but are suboptimal for equation-specific objectives such as fitting accuracy, which is caused by only using the cross-entropy loss between the generated sequence and the real sequence, we encode the expression symbol sequence and the constant sequence separately. For example, for the expression , we would have the skeleton . And constant sequence . It is worth noting that we replace constants with special symbols. The constants are encoded using a scientific-like notation where a constant C is represented as a tuple of the exponent and the mantissa , . Where belongs to integers between [-5,5] and belongs to [-1,1]. For example, for the constant 0.021 above, we use the placeholder C-1 in the skeleton, where is equal to 0.21 and is equal to -1, so .
@@ 3.1.3 Expression Decoder
The decoder autoregressively generates the symbols sequence (Skeleton) and the constants sequences given the encoder’s features. We can think of each token as a token in a text generation task. Then, n candidate expressions are obtained by sampling sequentially using beam search according to the predicted probability distribution of the next symbol. And we’ll get n constants sequences corresponding to n skeletons of expressions. Next, the constant placeholder ‘[C-5,…, C5]’ in the skeleton is replaced by the corresponding constant to obtain the expression with the rough constant. Finally, we fine-tune the rough constant with BFGS to obtain the optimal expression.
@@ 4.3 Ablation Study
To verify whether the introduction of contrastive learning plays a positive role in the algorithm performance, we conduct the ablation experiment as follows. On the premise of ensuring that the training data and other factors are the same, According to Equation 2, we did three sets of experiments, in which and were always equal to 1, and was set to 0.0, 0.1, and 1.0 in turn. The specific experimental results are shown in Table 1.
This can be clearly seen from the table. Other conditions being equal, as the proportion of contrastive learning loss () decreases gradually, the performance of the algorithm also decreases gradually. In addition, the overall performance with the contrastive loss ( = 1.0 and 0.1) is better than without the contrastive loss ( = 0.0). The contrastive loss makes the matching data features and skeleton features closer together in the feature space, and more importantly, it makes the non-matching data features and skeleton features further apart. In this way, the latter two features can be better fused.
@@ 4.4 Effects of modality alignment
To verify whether the extracted features before and after the introduction of contrastive learning are as expected: Data and skeleton features that match each other are close to each other in the feature space, whereas those that do not match are far away from each other. We first randomly generate 50 expression samples, feed these samples into the model without contrastive learning, and then obtain 50 pairs of features for the data and skeleton. We calculate the cosine distance between the extracted 50 data features and 50 skeleton features and then obtain the 50*50 cosine distance matrix A. Finally, we normalize the matrix A and visualize it to get fig 3(a). Similarly, we feed these samples into a model that uses contrastive learning, resulting in heatmap 3(b).
@@ 4.5 Impact of the size of the training data
To test the impact of pre-training data size on model test performance, we train our MMSR model on increasingly larger datasets. More specifically, we use datasets consisting of 10K, 100K, 1M, 5M, and 10M samples to train a model respectively. Except for the amount of data, every other aspect of training is the same as described above, and all are trained for the same epochs.
@@ 4.6 Effect of the introduction method of contrastive learning
Figure.3 shows the comparison plot between the One-step strategy and the Tow-step strategy. To prove the superiority of our adopted strategy, we conduct comparative experiments on all datasets. From figure 6, we can find that in terms of contrastive learning introduction method, the One-step strategy we adopted is slightly better than the Tow-step strategy in terms of effect. As the size of the dataset increases, this advantage becomes more and more obvious.
We believe that the specific reason for this phenomenon is that in the One-step strategy, contrastive loss and other losses are trained together. In the process of training, the feature extraction module and the feature fusion module are interrelated. This requires the feature extraction module not only to make the extracted data and skeleton features close in the feature space but also to make the extracted features better adapt to the later feature fusion module. So that the overall loss is smaller. In short, the one-step strategy can make each module better wearing in each other in the process of training together.
@@ Acknowledgements
Appendix A Supplementary material on the effect of data size on performance
This section shows the performance change trend of MMSR on Strogatz, R, Korns, Constant, Livermore, Black-box, Neat and Others datasets with the increase of training dataset size. From the figure, we can also see that the performance of MMSR increases with the growth of data size. This means that MMSR can theoretically get better results if we have more datasets.
Appendix B Supplementary material on one-step and two-step strategies
This section shows the performance trends of the one-step strategy and the two-step strategy on Strogatz, R, Korns, Constant, Livermore, Black-box, Neat and Others with the increase of the size of the training dataset. It can also be seen from the figure that with the growth of the data scale, no matter which strategy is used, the performance of the algorithm increases. However, the one-step strategy consistently outperforms the two-step strategy, and this trend becomes more and more obvious as the size of the data increases.
Appendix C Appendix: Test data in detail
Table 2,3,4 shows in detail the expression forms of the data set used in the experiment, as well as the sampling range and sampling number. Some specific presentation rules are described below
The variables contained in the regression task are represented as [].
signifies random points uniformly sampled between and for each input variable. Different random seeds are used for training and testing datasets.
indicates points evenly spaced between and for each input variable.
Name | Expression | Dataset |
Nguyen-1 | U | |
Nguyen-2 | U | |
Nguyen-3 | U | |
Nguyen-4 | U | |
Nguyen-5 | U | |
Nguyen-6 | U | |
Nguyen-7 | U | |
Nguyen-8 | U | |
Nguyen-9 | U | |
Nguyen-10 | U | |
Nguyen-11 | U | |
Nguyen-12 | U | |
Nguyen-2 | U | |
Nguyen-5 | U | |
Nguyen-8 | U | |
Nguyen-8 | U | |
Nguyen-1c | U | |
Nguyen-5c | ||
Nguyen-7c | U | |
Nguyen-8c | U | |
Nguyen-10c | U | |
Korns-1 | U | |
Korns-2 | U | |
Korns-3 | U | |
Korns-4 | U | |
Korns-5 | U | |
Korns-6 | U | |
Korns-7 | U | |
Korns-8 | U | |
Korns-9 | U | |
Korns-10 | U | |
Korns-11 | U | |
Korns-12 | U | |
Korns-13 | U | |
Korns-14 | U | |
Korns-15 | U | |
Jin-1 | U | |
Jin-2 | U | |
Jin-3 | U | |
Jin-4 | U | |
Jin-5 | U | |
Jin-6 | U |
Name | Expression | Dataset |
Neat-1 | U | |
Neat-2 | U | |
Neat-3 | U | |
Neat-4 | U | |
Neat-5 | U | |
Neat-6 | E | |
Neat-7 | E | |
Neat-8 | U | |
Neat-9 | E | |
Keijzer-1 | U | |
Keijzer-2 | U | |
Keijzer-3 | U | |
Keijzer-4 | U | |
Keijzer-5 | U | |
Keijzer-6 | U | |
Keijzer-7 | U | |
Keijzer-8 | U | |
Keijzer-9 | U | |
Keijzer-10 | U | |
Keijzer-11 | U | |
Keijzer-12 | U | |
Keijzer-13 | U | |
Keijzer-14 | U | |
Keijzer-15 | U | |
Livermore-1 | U | |
Livermore-2 | U | |
Livermore-3 | U | |
Livermore-4 | U | |
Livermore-5 | U | |
Livermore-6 | U | |
Livermore-7 | U | |
Livermore-8 | U | |
Livermore-9 | U | |
Livermore-10 | U | |
Livermore-11 | U | |
Livermore-12 | U | |
Livermore-13 | U | |
Livermore-14 | U | |
Livermore-15 | U | |
Livermore-16 | U | |
Livermore-17 | U | |
Livermore-18 | U | |
Livermore-19 | U | |
Livermore-20 | U | |
Livermore-21 | U | |
Livermore-22 | U |
Name | Expression | Dataset |
Vladislavleva-1 | U | |
Vladislavleva-2 | U | |
Vladislavleva-3 | U | |
Vladislavleva-4 | U | |
Vladislavleva-5 | U | |
Vladislavleva-6 | E | |
Vladislavleva-7 | E | |
Vladislavleva-8 | U | |
Test-2 | U | |
Const-Test-1 | U | |
GrammarVAE-1 | U | |
Sine | U | |
Nonic | U | |
Pagie-1 | E | |
Meier-3 | E | |
Meier-4 | ||
Poly-10 | E | |
Constant-1 | ||
Constant-2 | ||
Constant-3 | ||
Constant-4 | ||
Constant-5 | ||
Constant-6 | ||
Constant-7 | ||
Constant-8 | ||
R1 | ||
R2 | ||
R3 |
Appendix D Appendix: MMSR tests on AIFeynman dataset.
In our study, we conducted an evaluation of our novel symbol regression algorithm, termed MMSR, leveraging the AI Feynman dataset, which comprises a diverse array of problems spanning various subfields of physics and mathematics, including mechanics, thermodynamics, and electromagnetism. Originally, the dataset contains 100,000 data points; however, for a more rigorous assessment of MMSR’s efficacy, our analysis was deliberately confined to a subset of 100 data points. Through the application of MMSR for symbol regression on these selected data points, we meticulously calculated the values to compare the algorithm’s predictions against the true solutions.
The empirical results from our investigation unequivocally affirm that MMSR possesses an exceptional ability to discern the underlying mathematical expressions from a constrained sample size. Notably, the values achieved were above 0.99 for a predominant portion of the equations, underscoring the algorithm’s remarkable accuracy in fitting these expressions. These findings decisively position MMSR as a potent tool for addressing complex problems within the domains of physics and mathematics. The broader implications of our study suggest that MMSR holds considerable promise for a wide range of applications across different fields. Detailed experimental results are presented in Table 5 and Table 6.
Feynman | Equation | |
I.6.20a | 0.9999 | |
I.6.20 | 0.9991 | |
I.6.20b | 0.9881 | |
I.8.14 | 0.9024 | |
I.9.18 | 0.9926 | |
I.10.7 | 0.9872 | |
I.11.19 | 0.9999 | |
I.12.1 | 1.0 | |
I.12.2 | 1.0 | |
I.12.4 | 0.9999 | |
I.12.5 | 1.0 | |
I.12.11 | 0.9999 | |
I.13.4 | 0.9982 | |
I.13.12 | 1.0 | |
I.14.3 | 1.0 | |
I.14.4 | 0.9839 | |
I.15.3x | 0.9793 | |
I.15.3t | 0.9638 | |
I.15.10 | 0.9919 | |
I.16.6 | 0.9873 | |
I.18.4 | 0.9794 | |
I.18.12 | 0.9999 | |
I.18.16 | 0.9999 | |
I.24.6 | 0.9986 | |
I.25.13 | 1.0 | |
I.26.2 | 0.9999 | |
I.27.6 | 0.9895 | |
I.29.4 | 1.0 | |
I.29.16 | 0.9828 | |
I.30.3 | 0.9912 | |
I.30.5 | 0.9822 | |
I.32.5 | 0.9932 | |
I.32.17 | 0.9817 | |
I.34.8 | 1.0 | |
I.34.10 | 0.9917 | |
I.34.14 | 0.9992 | |
I.34.27 | 0.9999 | |
I.37.4 | 0.9827 | |
I.38.12 | 0.9999 | |
I.39.10 | 0.9999 | |
I.39.11 | 0.9914 | |
I.39.22 | 0.9976 | |
I.40.1 | 0.9816 | |
I.41.16 | 0.9213 | |
I.43.16 | 0.9981 | |
I.43.31 | 1.0 | |
I.43.43 | 0.9215 | |
I.44.4 | 0.8017 | |
I.47.23 | 0.9733 | |
I.48.20 | 0.8629 | |
I.50.26 | 0.9999 |
Feynman | Equation | |
II.2.42 | P | 0.8015 |
II.3.24 | 0.9813 | |
II.4.23 | 0.9972 | |
II.6.11 | 0.9883 | |
II.6.15a | 0.9221 | |
II.6.15b | 0.9917 | |
II.8.7 | 0.9816 | |
II.8.31 | 1.0 | |
II.10.9 | 0.9999 | |
II.11.3 | 0.9824 | |
II.11.7 | 0.8729 | |
II.11.20 | 0.7225 | |
II.11.27 | 0.9817 | |
II.11.28 | 0.9991 | |
II.13.17 | 0.9961 | |
II.13.23 | 0.9622 | |
II.13.34 | 0.9847 | |
II.15.4 | 0.9999 | |
II.15.5 | 0.9999 | |
II.21.32 | 0.9915 | |
II.24.17 | 0.9872 | |
II.27.16 | 0.9917 | |
II.27.18 | 0.9993 | |
II.34.2a | 0.9916 | |
II.34.2 | 0.9862 | |
II.34.11 | 0.9926 | |
II.34.29a | 0.9987 | |
II.34.29b | 0.8219 | |
II.35.18 | 0.9512 | |
II.35.21 | 0.8199 | |
II.36.38 | 0.9250 | |
II.37.1 | 0.9999 | |
II.38.3 | 0.9999 | |
II.38.14 | 0.9999 | |
III.4.32 | 0.9628 | |
III.4.33 | 0.9973 | |
III.7.38 | 0.9826 | |
III.8.54 | 0.9822 | |
III.9.52 | 0.7024 | |
III.10.19 | 0.9948 | |
III.12.43 | 0.9924 | |
III.13.18 | 0.9999 | |
III.14.14 | 0.9914 | |
III.15.12 | 0.9999 | |
III.15.14 | 0.9995 | |
III.15.27 | 0.9914 | |
III.17.37 | 0.9988 | |
III.19.51 | 0.9894 | |
III.21.20 | 0.7262 |